On a map, 1/3 in. represents 4 mi. If two cities are 19. in apart on the map, what is the actual distance between them?
Answer #1

1/3 inch. represents 4 mile

19. inch represent ?

\large \frac{19*4}{\frac{1}{3}}

\large \frac{19*4*3}{1}

\large {\color{Red} 228} miles.....